### v4

@@ 4 Sequence alignment of ART-simulated reads
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (4), and for this, we need to compute the SW distance. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
Distance bound, $d_{SW}\in\{1\%,2\%,5\%\}$: It is well known that short fragments frequently repeat in the genome and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled mappable ; reinventwillis . In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible. Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match. We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW distance between the pair ($v^{*}$ – see Eq. 4) and the optimal SW distance as a particular read length $Q$:
@@
where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW distance. We consider an alignment (with $Q=250$) to be successful if the resulting SW distance between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length. A $d_{SW}=2\%$ (default) is equivalent to a mismatch of around $4$ bases in a read of length $250$.
@@
Table 1: Performance of DNA-ESA with respect to baselines: The performance of DNA-ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig. 1. In addition, DNA-ESA is also compared with Bowtie-2 bowtie , a classical aligner. Comparisons are conducted across varying qualities (Phred score) of reads generated by ART huang_art_2012 . Additionally, an external PacBio CCS dataset of ($Q=250$ length-) reads from Chr. 2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and DNA-ESA. All baseline models utilize a dedicated vector store. For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations. Across models and ART-generated datasets, the performance of DNA-ESA supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$. With respect to reads from PacBio CCS, DNA-ESA performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW distance. Mean ($\mu$) and standard deviation ($\sigma$) across the top SW distance values ($v^{*}$) of reads successfully aligned in the case of DNA-ESA and Bowtie-2 models are reported.
4.1 Performance
@@
Table 1 reports the performance of DNA-ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec. 4), in addition to a direct comparison to DB2, mean and HD, max baselines (without any further finetuning), the best-performing baselines on pure reads identified in Sec. 3. In addition, we also compare DNA-ESA to Bowtie-2 bowtie , a conventional algorithmic aligner.
@@
Additionally, models are also evaluated on an external set of reads from PacBio CCN generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ). For (ii), the raw reads are $10$ kilobases long and for evaluation we consider random subset of $5000$ reads associated to Chr. 2333Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package minimap . Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of DNA-ESA ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
@@
We observe that: (A) DNA-ESA shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, DNA-ESA and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW distance). The results indicate that this constitutes a new state-of-the-art model for Transformer-based sequence alignment.
@@
Sweeping Top-K and the SW distance bound In Table 2, we report the performance of DNA-ESA across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$. Distance bound $d_{SW}<5\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments. Recall rates are comparable with different $d_{SW}$ values suggesting that when a read is successfully aligned, it is usually aligned to an objectively best match. Decreasing the top-$K$ per chromosome from $50\rightarrow 25$ does not substantially worsen performance ($<1\%$), indicating that the optimal retrievals are usually the closest in the embedding space. Additional ablation studies are presented in the Appendix.
@@
Table 2: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec. 4. DNA-ESA presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie. As expected, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
4.2 Task transfer from Chromosome 2
Are these results indicative of DNA-ESA’s adaptability to new genomic sequences rather than strict adherence to its training data? This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
[…]
Acc. $\uparrow$ F1 $\uparrow$ P $\uparrow$ R $\uparrow$
Train ART Reads 2 (seen in training) 3 (populated) Y (populated)
Chr. 2 $Q_{PH}\in[30,60]$ 99.74 $\pm$ 0.29 99.74 $\pm$ 0.29 $9$$\pm$ 0.01 99.49±0.46 98.87±0.59 99.99±0.01
Chr. 2 $Q_{PH}\in[60,90]$ 99.37 $\pm$ 0.49 99.36 $\pm$ 0.49 $9$$\pm$ 0.01 98.74±0.64 99.12±0.55 99.5±0.43
Table 3: DNA-ESA Performance on Chr. 2 and Task Transfer to Chr. 3, Y: DNA-ESA, trained on Chr. 2, is evaluated for aligning positive samples from Chr. 2 and negative samples from Chr. 3. Anticipated failure to align Chr. 3 reads (false positives) is due to significant SW distance. Metrics reported include weighted accuracy (Acc.), precision (P), F1-score (F1), and recall (R), for comparison with Table 1. [Grey] For Chr. 3 and Y, separate vector stores are populated using DNA-ESA representations learned using Chr. 2 {read,fragment} pairs. These results support the generalizing tendency of DNA-ESA to align sequences beyond the training reference.
Species $Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
Recall Top-K @50 $\uparrow$
Thermus Aquaticus [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
Acidobacteriota [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
Rattus Norwegicus [Chr. 1] 95.98±1.07 96.48±1.01
Rattus Norwegicus [Chr. 2] 97.99±0.78 96.49±1.01
Pan Troglodytes [Chr. 2A] 98.5±0.69 95.74±1.11
Pan Troglodytes [Chr. 2B] 95.99±1.08 96.25±1.05
Table 4: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr. 2, Testing on Diverse Species: DNA-ESA, trained on human Chr. 2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus. These species, are evaluated for read alignment recall using ART-generated reads. The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data.
Performance: Details on convergence are in Appendix 8. Table 3 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr. 2. This suggests DNA-ESA’s ability to generalize sequence alignment across chromosomes with different compositions. High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall. In Table 4, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (originstory, ). Even distantly related species like Thermus aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
5 Limitations and Future Work
DNA-ESA has the following challenges that we hope to overcome in future generations of the system: the computational speed, at 10K reads per minute, is a limiting factor for extensive genomic studies involving millions of reads. Despite this, the alignment task’s inherent parallelizability offers avenues for efficiency improvements. Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec. 9. We are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. Additionally, we aim to enhance DNA-ESA’s performance at shorter read lengths, exploring alternative training methods and integrating more diverse read/reference features.
6 Concluding Remarks
Current DNA sequence alignment methods, honed over decades, incorporate DNA-specific enhancements in both indexing and retrieval, relying on algorithmic approaches. However, we introduce an alternative data-driven paradigm, employing a Transformer-based DNA-ESA encoder. Our method performs sequence alignment through self-supervised learning, utilizing contrastive loss. While similar methods have previously found application in identifying approximate semantic similarity in written language, our approach surprisingly excels at identifying exact overlaps among DNA sequences. Empirical results underscore the model’s ability to embody the inherent structure of any DNA sequence, irrespective of location or species origin. A distinguishing feature of DNA-ESA is its ability to map both shorter reads and longer reference fragments into the same space by employing cosine similarity as a symmetric distance metric. This complements existing paradigms that have traditionally used a hierarchical and explicit comparison of base distributions.
DNA-ESA representations would find utility in other related tasks. If, for example, one had to align reads to the Pan Genome liao instead of a single reference, one can populate a single vector store with fragments from all $47$ genome assemblies. Then given a read, the nearest fragment to the read can belong to any reference variant. Note that in this case, we avoid having to construct a separate index for each of the references as is usually done in conventional methods.
One can similarly repurpose the representation space – originally intended for Sequence Alignment – to address the complementary task of Genome Assembly. We conclude this paper with a setup that provides initial proofpoints under reasonable simplifying assumptions about read lengths and overlaps among the reads. The resulting framework will be formalized and benchmarked with reads from actual experiments in our future work.
6.1 A setup for de-novo Genome Assembly using DNA-ESA
Consider a set of reads, $R:=\{r_{1},r_{2},\dots,r_{C}\}$ from which we would like to estimate a reference $\mathcal{R}$. The de-novo assembly task is to estimate an ordering of the reads in $R$, $\hat{R}:=(r_{i},\dots,r_{j},\dots,r_{k})$ such that a concatenation (after removing overlaps) is approximately the original reference. In typical setups, all pairs of reads would need to be compared in order to find those that overlap and construct longer chains of the assembly in an incremental fashion. This may be visualized as a complete graph $K_{C}$ with $C$ nodes (reads) and $C^{2}$ edges (constraints). To avoid the quadratic complexity, modern techniques such as Overlap-Layout-Consensus (OLC) or de Bruijn Graph (DBG) approaches assembly have adopted hot-start heuristics such as base overlap and shared K-mer counts to reduce the number of pairwise read comparisons.
The representation space modeled by DNA-ESA admits a comparable assembly setup. Fig. 2 suggested emergent 1D manifolds in the representation space – comprising the embeddings of the reads – $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ – such that the relative positions of fragments along the manifold correlated to the relative locations of the fragments along the reference. Therefore, a pairwise read-read comparison within a smaller radius in the embedding space would be sufficient to find reads that overlap. And more generally, a walk along the 1D manifold would constitute a near-optimal assembly.
Figure 5: Motivating the task of de-novo assembly using the distance adjacency matrix of the UMAP-generated $k$-nearest neighbor ($k=10$) network of reads from Thermus Aquaticus: We assume that reads $r_{1},r_{2},\dots,r_{C}$ belonging to a species are embedded into the DNA-ESA embedding space $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ using a model trained on a known reference genome belonging to a different species. For this illustrative numerical example we ensure that reads are created with the following properties: first, no read is a substring of any other read (efficient algorithms for the case where reads can be substrings of others will be covered in our future work); second every read has a corresponding read that it overlaps with for each end, and finally, the union of the reads cover the ground-truth reference genome. Thus, for every read $r_{i}$ there exist reads $r_{k}$ and $r_{j}$ such that the ordered triplet $(r_{j},r_{i},r_{k})$ constitutes a fragment of the reference. In a perfect embedding scenario, the closest neighbors of $r_{i}$ (i.e. $k=1$ in an $k$-NN search) in the embedding space should be $r_{j}$ ad $r_{k}$. For this ideal case, if a network is constructed where each read (node) is connected to its two nearest neighbors, and the nodes are arranged based on the order in which they appear in the reference chromosome, one would get a bi-diagonal adjacency matrix. When ground truth is not known, the adjacency matrix will be a permuted version of a bi-diagonal matrix, but a complete assembly can still be performed by a greedy algorithm that starts with any node and pieces together the neighboring reads in the embedded space in time that grows linearly in the number of reads. Simulated reads, each of length $500$ and with random overlaps of lengths drawn uniformly $\mathcal{U}([150,250])$ with its neighboring reads at each end, are sampled from $4$ of the chromosomes in Thermus Aquaticus, and are embedded into the DNA-ESA embedding space trained with a reference human genome. The sorted adjacency matrix of the UMAP umap -generated $k$-NN ($k=10$) network is presented above for each chromosome. The dark off-diagonal distance values in the $k$-NN adjacency graph with $k=10$ show that the distance properties of actual embeddings of simulated reads from Thermus Aquaticus are almost coincident with that of the ideal case discussed above. Moreover, just as in the ideal case, by finding an approximate solution to the Traveling Salesman Problem in the unsorted network, one can find a walk (the assembly) that is close to the original reference. The results are presented in Table 5.
We provide an initial computational implementation of this idea: (a) The k-nearest neighbor (kNN) graph $G(R,d)$, representing the distances between close-by reads, is computed using the density-aware corrections provided by the UMAP library umap : graph adjacency matrices for each of the chromosomes is visualized in Fig. 5. For this figure, the several reads are ordered manually as they appear in the assembly in order to highlight the connectivity pattern; the bright off-diagonal values indicate a chain-like structure corresponding to the observed manifold; (b) A Hamiltonian cycle computed through the graph that has the lowest total distance corresponds to an assembly, $\hat{R}$. This involves solving the Traveling Salesman Problem (TSP) – we use the 3/2-approximate Christofides algorithm christofides . After a walk across the reads is generated, a second pass through the walk greedily concatenates adjacent reads $(r_{i},r_{i+1})$ by computing the alignment using the SW distance and removing duplicates and overlaps. In cases where adjacent reads do not have any alignment to one another – the walk isn’t guaranteed to be perfect since the DNA-ESA model is heuristical and the TSP algorithm is approximate – the current assembly part is stowed away as a contig contig , and a new contig is instantiated with the incoming read. The fewer the contigs, the better the manifold in the representation space correlates to the ordering of the true reference.
Assembly is conducted across $4$ of the shorter chromosomes in the Thermus Aquaticus genome. We consider (long) reads of length $Q=500$, which have no indels, and the entire corpus of reads is ensured to have full coverage across each chromosome. Each read has an overlap of $\mathcal{U}([150,250])$ bases with another read in the corpus of reads. QUAST quast metrics are reported in Table 5.
Size (in kb) $\uparrow$ No. of Contigs $\downarrow$ N50 (in kb) $\uparrow$ NGA50 (in kb) $\uparrow$
Thermus Aquaticus NZ_CP010823.1 14.4 1 14.4 14.4
NZ_CP010824.1 16.6 1 16.6 16.6
NZ_CP010825.1 69.9 22 6.0 4.4
NZ_CP010826.1 78.7 15 10.3 10.3
Table 5: DNA-ESA performance on de-novo assembly of the Thermus Aquaticus genome: The solution to the approximate TSP applied on the DNA-ESA read-read network generates a set of contigs that can be evaluated with respect to the original reference using QUAST quast . For the two shorter chrosomomes, the resulting assembly was perfect; the entire reference is encapsulated by one contig. Even in the case with longer chromosomes, the number of contigs are still few and of significant length (as indicated by the high N(G)50 score) suggesting a strong correlation between the distance metric in the embedding space and the ordering of the corresponding reads in the assembly.
N50 refers to the length of a contig (in bases) such that the contigs that exceed this length span 50% of the reference genome. NGA50 is a reference-aware N50 metric. The fewer contigs, large N(GA)50 scores in comparison to the large size of the chromosomes, when considered together, demonstrate that the reads are indeed ordered in a relative position-aware manner in the DNA-ESA embedding space with respect to a reference making the de-novo assembly task viable.
{appendices}
7 Appendix
8 A. Convergence
Fig. 6 plot the convergence of the DNA-ESA encoder model discussed in the main text. We see convergence after $\sim$20k steps.
((a)) Trained on the Human Genome
((b)) Trained on the Human Chromosome 2
Figure 6: DNA-ESA convergence plots. Both plots show the loss pr. step (grey). For clarity, we smooth the loss using a moving average (black).
9 B. Complexity of computing alignment
9.1 Cost of constructing a new representation
Computing the embedding $\mathcal{E}$ of a sequence of length $F$ using DNA-ESA encoding – a typical Transformer-based attention architecture – has the following computation complexity:
$\mathcal{O}(LH*(F^{2}*d+d^{2}*F))\Rightarrow\mathcal{O}(F^{2}*d+d^{2}*F)$ (8)
Where $d$ is the embedding dimension of the model, $L$ is the number of layers in the Transformer and $H$ is the number of heads per layer. As $d$ is a controllable parameter for the model, we can further simplify:
$\mathcal{O}(F^{2}*d+d^{2}*F)\Rightarrow\mathcal{O}(F^{2})$ (9)
The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (beltagy_longformer_2020, ; kitaev_reformer_2020, ) have developed shortcuts to reduce the cost. These have already been applied to DNA sequence modeling  (nguyen_hyenadna_2023, ).
9.2 Vector Store Upstream
Vector store $\mathcal{D}$ is populated once (in bulk) with encoded fragment-length sequences drawn from the entire genome; constant time complexity $C$ to upload $<10M$ vectors.
9.3 Retrieval Cost
Given a new embedding, $\epsilon(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome. In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, $\epsilon$ scales logarithmically with $G$. This is indeed the key benefit of using such databases.
9.4 Fine-grained Alignment
Existing libraries/algorithms (e.g. the Smith–Waterman algorithm) can identify the alignment between a fragment sequence (of length $F$) and read (of length $Q$) in $\mathcal{O}(FQ)$.
9.5 Total Complexity
Total complexity involves (a) constructing the representation of a read; (b) querying the vector store; (c) running fine-grained alignment with respect to the $K$ returned reference fragment sequences:
$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))$
$\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
10 C. Ablation studies
10.1 Without diversity priors in top-K
In Table 6, we report the performance without diversity priors used in the retrieval step: i.e. the nearest-K neighbors in the embedding space are selected from the entire set of fragments spanning the genome rather than uniformly sampling from each chromosome. The performance predicably falls in comparison to those reported in Table 1 since fewer fragments scattered unevenly across the different chromosomes are being retrieved per read.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
[…]
0.0 0.0 98.56 $\pm$ 0.67 97.60 $\pm$ 0.85 98.56 $\pm$ 0.67 97.76 $\pm$ 0.82 97.76 $\pm$ 0.82 97.44 $\pm$ 0.88
0.01 0.01 98.8 $\pm$ 0.62 98.4 $\pm$ 0.71 98.8 $\pm$ 0.62 99.2 $\pm$ 0.52 98.4 $\pm$ 0.71 98.4 $\pm$ 0.71
Top-50 global
0.0 0.0 91.6 $\pm$ 1.49 92.4 $\pm$ 1.43 93.8 $\pm$ 1.31 90.4 $\pm$ 1.58 93.8 $\pm$ 1.31 94.8 $\pm$ 1.21
0.01 0.01 89.6 $\pm$ 1.64 91.6 $\pm$ 1.49 96 $\pm$ 1.07 94.4 $\pm$ 1.25 94.4 $\pm$ 1.25 93.2 $\pm$ 1.36
Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec. 4. When compared to the main results presented in Table 1, it is evident that the addition of diversity priors results in an improvement of $\sim$ in recall. Similar to the result presented in the main text, performance improves with larger search radius in the vector store (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
10.2 Across read lengths
In Table 7, the performance of DNA-ESA is reported across read lengths. We observe Zero-shot performance at longer read lengths: The model performs better at longer read lengths (even exceeding the read length bound established during training $\mathcal{U}[150,500]$); while evaluating for longer reads, we make sure to guarantee that the reads exist as subsequences of fragments. Improving the performance at shorter read lengths is the subject of future work.
I D $Q=200$ $Q=225$ $Q=250$
$Q_{PH}\in[30,60]$
0 0 99 $\pm$ 0.57 98.6 $\pm$ 0.67 99 $\pm$ 0.57
0.01 0.01 97.4 $\pm$ 0.88 98.8 $\pm$ 0.62 98.6 $\pm$ 0.67
$Q_{PH}\in[60,90]$
0 0 98.6 $\pm$ 0.67 98.6 $\pm$ 0.67 98.9 $\pm$ 0.59
0.01 0.01 98.6 $\pm$ 0.67 99 $\pm$ 0.57 98.6 $\pm$ 0.67
Table 7: DNA-ESA recall performance across read lengths: Performance of DNA-ESA is higher for longer reads including those lengths on which the model was not trained ($Q\in\mathcal{U}([150,500])$). Shorter reads are more challenging for the model potentially due to replicates found across the reference.
11 Competing interests
No competing interest is declared.
12 Author contributions statement
VR devised the task and framework and assisted with writing. MP aided in writing and bridging NLP and Bioinformatics communities. LSB contributed to writing and brainstorming. PH and KCE realized the framework and assisted with writing. SR ran several experiments to improve model performance. LM and TG reviewed existing conventional baselines.

### v5

@@ Abstract
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (6), and for this, we need to compute the SW score. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
Distance bound, $d_{SW}\in\{1\%,2\%,5\%\}$: It is well known that short fragments frequently repeat in the genome and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled mappable ; reinventwillis . In this case, $q^{*}\neq\hat{q}^{*}$, but the SW score is the minimum possible. Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match. We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW score between the pair ($v^{*}$ – see Eq. 6) and the optimal SW score as a particular read length $Q$:
@@
where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW score. We consider an alignment (with $Q=250$) to be successful if the resulting SW score between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length. A $d_{SW}=2\%$ (default) is equivalent to a mismatch of around $4$ bases in a read of length $250$.
@@
Table 1: Performance of RDE with respect to baselines: The performance of RDE on ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig. 1. In addition, RDE is also compared with Bowtie-2 bowtie , a classical aligner. Comparisons are conducted across varying qualities (Phred score) of reads generated by ART huang_art_2012 . Additionally, an external PacBio CCS dataset of ($Q=250$ length-) reads from Chr. 2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and RDE. All baseline models utilize a dedicated vector store. For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations. Across models and ART-generated datasets, the performance of RDE supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$. With respect to reads from PacBio CCS, RDE performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW score. Mean ($\mu$) and standard deviation ($\sigma$) across the top SW scores($v^{*}$) of reads successfully aligned in the case of RDE and Bowtie-2 models are reported.
4.1 4.1 Performance
@@
Table 1 reports the performance of RDE on ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec. 4), in addition to a direct comparison to DB2, mean and HD, max baselines (without any further finetuning), the best-performing baselines on pure reads identified in Sec. 3. In addition, we also compare RDE to Bowtie-2 bowtie , a conventional algorithmic aligner.
@@
Additionally, models are also evaluated on an external set of reads from PacBio CCS generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ). For (ii), the raw reads are $10$ kilobases long and for evaluation we consider random subset of $5000$ reads associated to Chr. 2. Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of RDE ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
@@
We observe that: (A) RDE shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, RDE and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW score). The results indicate that this constitutes a new state-of-the-art model for Transformer-based sequence alignment.
@@
Sweeping Top-K and the SW distance bound In Table 2, we report the performance of RDE across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$. Distance bound $d_{SW}<5\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments. Recall rates are comparable with different $d_{SW}$ values suggesting that when a read is successfully aligned, it is usually aligned to an objectively best match. Decreasing the top-$K$ per chromosome from $50\rightarrow 25$ does not substantially worsen performance ($<1\%$), indicating that the optimal retrievals are usually the closest in the embedding space. Several additional experiments are presented in the SI.
@@
Table 2: Sequence alignment recall of RDE sweeping top-K and $d_{SW}$: The various parameters are described in Sec. 4. RDE presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie. As expected, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
5 5. Limitations and Future Work
While our first-generation reference-free DNA embedding (RDE) models provide alignment performance that almost matches the performance of traditional aligners (refined over decades), there is room for considerable improvement. First, our RDE models have been trained using samples that span only around 2% of the human genome as shown in the Sec-A of the Supporting Material(SM). One needs to explore if the embedding properties of our models can be further improved by diversifying and augmenting training data and optimizing over the model parameters. The validation of RDE models can be performed only through various genomic tasks. In terms of the task of alignment, we plan to explore the following: (i) The current off-the-shelf implementation has a speed of around 10K reads per minute; while the existing aligners can achieve close to 1M reads per minute by trading off alignment accuracy. As detailed in SM Sec. B, we are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. We believe ESA can be made much faster; (ii) Optimization of the training of the RDE models to improve performance for shorter reads. As shown in see Table 2, Page 3 of SM the existing model performs well on longer reads; we want to develop RDE models that specialize to aligning shorter reads; and (iii) quantifying how embeddings change as subsequences are edited; this will help in better alignment accuracy. It will also help in correlating the geometry of the embedding space to the structure of DNA sequences.
6 6. Concluding Remarks
We have introduced a novel Reference-Free DNA Embedding (RDE) framework: it creates an embedding of any given DNA subsequence irrespective of the reference genome from which it is sampled. Such models capture similarity among genome subsequences in a purely data-driven manner (without explicitly computing edit-distance measures) and embeds DNA sequences of differing lengths into the same embedding space to facilitate search and assembly. Once such RDE models are constructed their expressive power and limitations can be assessed by performing multiple challenging genomic tasks in a flexible manner across different species.
As a first validation and application of RDE, we developed an Embed-Align-Search (ESA) framework where alignment of reads to a reference genome is achieved by a near-neighbor search in the embedding space: only the fragments of reference genome with embeddings nearest to the embedding of the given read need to be considered. Thus, a global search in a giga-bases long reference genome sequence is reduced to a local search in a vector space. Detailed experiments showed that the performance of even this first-generation RDE methodology is comparable to that of traditional aligners, which have been refined over decades. Future differentiating application of RDE to the alignment task might lie in aligning reads to graph- or pan-genomes liao , where the reference genome has multiple variants. Indexing such pangenomes is a difficult task in traditional aligners; however, in ESA one can store the RDE embeddings of the different fragments in the pangenome, and then perform a near-neighbor search in the embedding space. Thus, the ESA methodology does not have to be significantly altered, and alignment of reads to pangenomes constitutes an important future task.
Going forward we aim to answer the following important questions:
  • How reference-free is our RDE model? To benchmark the generalizability of our RDE model we have performed several cross-species alignment tasks as summarized in Tables 3 and 4 in the SM. In particular, we have used an RDE trained using human genome to align reads to reference genomes belonging to a wide range of species. The results are encouraging and the design of universal RDEs –that would generate effective embeddings of subsequences sampled from the genome of every species – is another direction of future research.
  • Is the RDE capable of generalizing to other tasks like de novo Genome Assembly? One can also harness the RDE space to address the complementary task of de novo Genome Assembly, where given a set of reads one assembles a genome by stitching together the reads, without using any reference. Genome assembly is considered to be a much more challenging task than alignment. In SM Sec. E, we provide initial proofpoints of a genome assembly framework, under reasonable simplifying assumptions about read lengths and overlaps among the reads. We show how the RDE model – trained on exemplars from human genome – can effectively embed simulated reads from a very different species, Thermus Aquaticus. We find that the read embeddings form a low-dimensional manifold, which can be efficiently searched locally to obtain high-fidelity genome assembly from the simulated reads. Our future work on genome assembly will focus on improving RDE models to obtain better embeddings, and also on developing local-search based algorithms that can accurately map low-dimensional manifolds (comprised by the embeddings of reads) into linear DNA sequences.